# Continuous Temporal Domain Generalization

### v1

@@ Abstract
Temporal Domain Generalization (TDG) addresses the challenge of training predictive models under temporally varying data distributions. Traditional TDG approaches typically focus on domain data collected at fixed, discrete time intervals, which limits their capability to capture the inherent dynamics within continuous-evolving and irregularly-observed temporal domains. To overcome this, this work formalizes the concept of Continuous Temporal Domain Generalization (CTDG), where domain data are derived from continuous times and are collected at arbitrary times. CTDG tackles critical challenges including: 1) Characterizing the continuous dynamics of both data and models, 2) Learning complex high-dimensional nonlinear dynamics, and 3) Optimizing and controlling the generalization across continuous temporal domains. To address them, we propose a Koopman operator-driven continuous temporal domain generalization (Koodos) framework. We formulate the problem within a continuous dynamic system and leverage the Koopman theory to learn the underlying dynamics; the framework is further enhanced with a comprehensive optimization strategy equipped with analysis and control driven by prior knowledge of the dynamics patterns. Extensive experiments demonstrate the effectiveness and efficiency of our approach.
@@ 1 Introduction
In practice, the distribution of training data often differs from that of test data, leading to a failure in generalizing models outside their training environments. Domain Generalization (DG) [43; 40; 17; 10; 34] is a machine learning strategy designed to learn a generalized model that performs well on the unseen target domain. The task becomes particularly pronounced in dynamic environments where the statistical properties of the target domains change over time [12; 28], prompting the development of Temporal Domain Generalization (TDG) [22; 33; 35; 1; 49; 50]. TDG recognizes that domain shifts are temporally correlated. It extends DG approaches by modeling domains as a sequence rather than as categorical entities, making it especially beneficial in fields where data is inherently time-varying.
@@
However, in practice, data may not always occur or be observed at discrete, regularly spaced time points. Instead, events and observations unfold irregularly and unpredictably in the time dimension, leading to temporal domains distributed irregularly and sparsely over continuous time. Formally, this paper introduces a problem termed Continuous Temporal Domain Generalization (CTDG), where both seen and unseen tasks reside in different domains at continuous, irregular time points. Fig. 1 illustrates an example of public opinion prediction after political events via Twitter data. Unlike the assumption in traditional TDG that the temporal regularly domains, the data are collected for the times near political events that may occur in arbitrary times. In the meanwhile, the domain data evolve constantly and continuously over time, e.g., active users increase, new friendships are formed, and age and gender distribution changes. Correspondingly, the ideal classifier should gradually change with the domain at random moments to counter the data distribution change over time. Finally, we are concerned with the state of the predictive model at any moment in the future. CTDG is ubiquitous in other fields. For example, in disaster management, relevant data are collected during and following disasters, which may occur at any time throughout the year. In healthcare, critical information about diagnosis and treatment is typically only documented during episodes of care rather than evenly throughout the lifetime. Hence, CTDG task necessitates the characterization of the continuous dynamics of irregular time-distributed domains, which cannot be handled by existing TDG methods designed for discrete-dynamics and fixed-interval times.
@@
To address all the challenges, we propose the Koopman operator-driven continuous temporal domain generalization framework (Koodos). Specifically, the Koodos framework articulates the evolutionary continuity of the predictive models in CTDG, then leverages a continuous dynamic approach to model its smooth evolution over time. Koodos further simplifies the nonlinear model system by projecting them into a linearized space via the Koopman Theory. Finally, Koodos provides an interface that reveals the internal model dynamic characteristics, as well as incorporates priori knowledge and constraints directly into the joint learning process.
2 Related Work
@@
Domain Generalization (DG) and Domain Adaptation (DA). DG approaches attempt to learn a model from multiple source domains that generalize well to an unseen domain [32; 31; 23; 2; 9]. Existing DG methods can be classified into three strategies [43]: (1) Data manipulation techniques, such as data augmentation [39; 40; 50] and data generation [27; 34]; (2) Representation learning focuses on extracting domain-invariant features [13; 14] and disentangling domain-shared from domain-specific features [25]; (3) Learning strategies encompass ensemble learning [29], meta-learning [21; 10], and gradient-based approaches [17]. Unlike DG, DA methods require simultaneously accessing source and target domain data to facilitate alignment and adaptation [44; 45; 24]. The technique includes domain-invariant learning [13; 41; 42; 47], domain mapping [3; 16; 11; 26], ensemble methods [37], and so on. Both DG and DA are limited to considering generalization across categorical domains, which treats domains as individuals but ignores the smooth evolution of them over time.
@@
Temporal Domain Generalization (TDG). TDG is an emerging field that extends traditional DG techniques to address challenges associated with time-varying data distributions. TDG decouples time from the domain and constructs domain sequences to capture its evolutionary relationships. S-MLDG [22] pioneers a sequential domain DG framework based on meta-learning. Gradient Interpolation (GI) [33] proposes to extrapolate the generalized model by supervising the first-order Taylor expansion of the learned function. LSSAE [35] deploys a probabilistic framework to explore the underlying structure in the latent space of predictive models. DRAIN [1] constructs a recurrent neural network that dynamically generates model parameters to adapt to changing domains. TKNets [50] minimize the divergence between forecasted and actual domain distributions to capture temporal patterns.
@@
Despite these studies, traditional TDG methods are limited to requiring domains presented in discrete time, which disrupts the inherent continuity of changes in data distribution, and the generalization can only be carried forward by limited steps. No work treats time as a continuous variable, thereby failing to capture the full dynamics of evolving domains and generalize to any moment in the future.
@@
Continuous Dynamical Systems (CDS). CDS are fundamental in understanding how systems evolve without the constraints of discrete intervals. They are the study of the dynamics for systems defined by differential equations. The linear multistep method or the Runge-Kutta method can solve the Order Differential Equations (ODEs) [15]. Distinguishing from traditional methods, Neural ODEs [7] represent a significant advancement in the field of CDS. It defines a hidden state as a solution to the ODEs initial-value problem, and parameterizes the derivatives of the hidden state using a neural network. The hidden state can then be evaluated at any desired time using a numerical ODEs solver. Many recent studies have proposed on which to learn differential equations from data [36; 18; 19; 30].
@@ 3 Problem Definition
Unlike traditional temporal domain generalization approaches [33; 1; 46; 50] that divide time into discrete intervals and require domain data be collected at fixed time steps, the CTDG problem treats time as a continuous variable, allowing for the time points in training and test set to be any arbitrary positive real numbers, and requiring models to deal with the temporal domains as continuous streams. CTDG problem poses several unprecedented, substantial challenges in: 1) Characterizing the continuous dynamics of data and how that decides model dynamics; 2) Modeling the low-dimensional continuous dynamics of the predictive model, which are embedded in high-dimensional space due to over-parametrization; and 3) Optimizing and analyzing the continuous predictive model system, including the application of inductive biases to control its behavior.
@@ 4 Methodology
Figure 2: Macro-flows and Micro-constraints in Proposed Model Framework
@@ 4.1 Characterizing the Continuous Dynamics of the Data and the Model
Here, $\mathcal{L}$ symbolizes the loss function tailored to the prediction task. By employing Eq. 7, the model dynamics are defined for all observation time, and $\phi$ can be obtained by optimizing Eq. 5 and Eq. 6 via gradient descent. After that, the parameters of the predictive model can be calculated at any specific time using ODEs solvers.
@@ 4.2 Modeling Nonlinear Model Dynamics by Koopman Operators
The aforementioned learning approaches are, however, limited in efficient modeling and prediction due to the entangled nonlinear dynamics in the high-dimensional parameters space. Identifying strongly nonlinear dynamics through such space is particularly susceptible to overfitting [5], leading to future predictive models eventually diverging from bifurcation to chaos. Despite the complexity of $h$, it is governed by the data dynamics $f$, which are typically low-dimensional and exhibit simple, predictable patterns. This suggests that the governing equations for $h$ are sparse within the over-parameterized space, and thus allow for a more simplified and manageable representation in a properly transformed space. Motivated by this, we propose leveraging the well-established Koopman Theory [20; 4] to represent these complex parameter dynamics. Koopman Theory provides a method for global linearization of any nonlinear dynamics. It expresses the complex dynamic system as an infinite-dimensional Koopman operator acting on the Hilbert space of the system state measurement functions, in which the nonlinear dynamics will become linearized.
@@ 4.3 Joint Optimization of Model and Its Dynamics with Prior Knowledge
We introduce a comprehensive optimization approach designed to ensure that the system accurately captures the dynamics of the data. This process requires the joint optimization of several interconnected components under the optional constraint of additional priori knowledge about the underlying dynamics: the predictive model parameters $\theta_{1:T}$, the transformation functions $\varphi$ and $\varphi^{-1}$, and the Koopman operator $\mathcal{K}$. Our primary objectives are threefold: 1) Ensuring high prediction accuracy, 2) Maintaining consistency of parameters across different representations and transformations, and 3) Learning the Koopman invariant subspaces effectively. Fig. 2(b) illustrates the role of each constraint within our system: we manage two sets of states, intrinsic and integral, aligning across three spaces.
Predictive Model Parameters ($\theta$): Each observation time corresponds to a predictive model, which is tasked with generating predictions and serving as the initial values for the dynamical system. They are primarily optimized to minimize the prediction error $L_{pred}$ of different domains:
$L_{pred}=\sum\nolimits_{i=1}^{T}\mathcal{L}(Y_{i},g(X_{i};\theta_{i})).$ (10)
Koopman Network Parameters ($\varphi$, $\varphi^{-1}$, $\mathcal{K}$): We estimate a function that transforms data into Koopman space. Meanwhile, it is vital that the parameters remain stable and consistent when transformed between different representations. The targets are realized by three loss constraints:
@@
    Reconstruction Loss: An auto-encoder is leveraged to map parameter space to Koopman space by encoder $\varphi$ and back to the original space by decoder $\varphi^{-1}$. $L_{recon}$ ensures consistency between $\theta$ and its reconstructed form via transformations:
@@
    Dynamic Fidelity Loss: This term ensures that the transformation produces a Koopman invariant subspaces in which the dynamics can forward correctly. It measures the fit of the Koopman operator $\mathcal{K}$ against transferred parameter:
@@
    Consistency Loss: It measures the consistency between the original and dynamic parameters in the model space:
@@
Additionally, we reload the dynamically integral parameters back into the predictive model to enhance its predictive capability, quantified by $L_{pred2}=\sum_{i=1}^{T}\sum_{j=1}^{i}\mathcal{L}(Y_{i},g(X_{i};\theta^{j\rightarrow i}_{i}))$. Finally, the system optimizes the following combined loss to refine all components simultaneously:
$\{\theta_{1:T},\varphi,\varphi^{-1},\mathcal{K}\}=\mathop{\arg\!min}_{\theta_{1:T},\varphi,\varphi^{-1},\mathcal{K}}(\alpha L_{pred}+\alpha L_{pred2}+\beta L_{recon}+\gamma L_{dyna}+\beta L_{consis}),$ (14)
with $\alpha$, $\beta$, and $\gamma$ as tunable weights balancing the loss magnitude of each space.
During inference, given the moment $t_{s}$, the model uses the state from the nearest observation moment $t_{obs}$ as an initial value, integrating over the interval $[t_{obs},t_{s}]$ on Koopman space to estimate the predictive model state at the desired test moment.
Analysis and Control of Predictive Models. Integrating Koopman theory into continuous temporal domain modeling facilitates the application of optimization, estimation, and control techniques, particularly through the spectral properties of the Koopman operator. We remark that the Koopman operator serves as a pivotal interface for analyzing and controlling the dynamics of predictive models. Constraints imposed on the Koopman space will be equivalently mapped to the parameter space. For instance, the eigenvalues of $\mathcal{K}$ are crucial as they provide insights into the system’s stability and dynamics, as illustrated below:
@@
    System Assessment. A predictive model is considered stable if all $\lambda_{i}$ have negative real parts. Conversely, the presence of $\lambda_{i}$ with a positive real part indicates instability. $\lambda_{i}$ with zero real parts suggest potential oscillatory behavior. By examining the locations of poles in the complex plane, the long-term behavior of the predictive model can be assessed, which helps us identify whether the generalized model is likely to collapse in the future.
@@
    Behavioral Constraints. Adding explicit constraints to $\mathcal{K}$ can guide the system toward desired behaviors. This process not only facilitates the incorporation of priori knowledge about the data but also tailors the system to specific functional requirements. To name just a few, if the data dynamics are known to be periodic, configuring $\mathcal{K}$ as $\mathcal{K}=B-B^{T}$, with $B$ as learnable parameters, ensures that the model system exhibits consistent periodic behavior since the eigenvalues of $B-B^{T}$ are purely imaginary values. Additionally, employing a low-rank approximation such as $\mathcal{K}=UV^{T}$, with $U,V\in\mathbb{R}^{n\times k}$ and $k<n$, allows the model to concentrate on the most significant dynamical features and explicitly control the degrees of freedom of the predictive model dynamics.
@@
Theorem 2 (Superiority of CTDG over discrete-time methods (informal)).
A continuous-time method has a smaller or equal error compared to a discrete-time method for approximating temporal distribution drift, due to its consideration of the integral of changes over time.
A detailed model complexity analysis and the formal version and proof of Theorem 2 are provided in Appendix A.1.4 and Appendix C.

### v2

@@ Abstract
Temporal Domain Generalization (TDG) addresses the challenge of training predictive models under temporally varying data distributions. Traditional TDG approaches typically focus on domain data collected at fixed, discrete time intervals, which limits their capability to capture the inherent dynamics within continuous-evolving and irregularly-observed temporal domains. To overcome this, this work formalizes the concept of Continuous Temporal Domain Generalization (CTDG), where domain data are derived from continuous times and are collected at arbitrary times. CTDG tackles critical challenges including: 1) Characterizing the continuous dynamics of both data and models, 2) Learning complex high-dimensional nonlinear dynamics, and 3) Optimizing and controlling the generalization across continuous temporal domains. To address them, we propose a Koopman operator-driven continuous temporal domain generalization (Koodos) framework. We formulate the problem within a continuous dynamic system and leverage the Koopman theory to learn the underlying dynamics; the framework is further enhanced with a comprehensive optimization strategy equipped with analysis and control driven by prior knowledge of the dynamics patterns. Extensive experiments demonstrate the effectiveness and efficiency of our approach. The code can be found at: https://github.com/Zekun-Cai/Koodos.
@@ 1 Introduction
In practice, the distribution of training data often differs from that of test data, leading to a failure in generalizing models outside their training environments. Domain Generalization (DG) [49; 46; 21; 14; 40] is a machine learning strategy designed to learn a generalized model that performs well on the unseen target domain. The task becomes particularly pronounced in dynamic environments where the statistical properties of the target domains change over time [16; 33; 4], prompting the development of Temporal Domain Generalization (TDG) [27; 39; 41; 2; 55; 57]. TDG recognizes that domain shifts are temporally correlated. It extends DG approaches by modeling domains as a sequence rather than as categorical entities, making it especially beneficial in fields where data is inherently time-varying.
@@
However, in practice, data may not always occur or be observed at discrete, regularly spaced time points. Instead, events and observations unfold irregularly and unpredictably in the time dimension, leading to temporal domains distributed irregularly and sparsely over continuous time. Formally, this paper introduces a problem termed Continuous Temporal Domain Generalization (CTDG), where both seen and unseen tasks reside in different domains at continuous, irregular time points. Fig. 1 illustrates an example of public opinion prediction after political events via Twitter data. Unlike the assumption in traditional TDG that the temporal regularly domains, the data are collected for the times near political events that may occur in arbitrary times. In the meanwhile, the domain data evolve constantly and continuously over time, e.g., active users increase, new friendships are formed, and age and gender distribution changes. Correspondingly, the ideal classifier should gradually change with the domain at random moments to counter the data distribution change over time. Finally, we are concerned with the state of the predictive model at any moment in the future. CTDG is ubiquitous in other fields. For example, in disaster management, relevant data are collected during and following disasters, which may occur at any time throughout the year. In healthcare, critical information about diagnosis and treatment is typically only documented during episodes of care rather than evenly throughout the lifetime. Hence, the CTDG task necessitates the characterization of the continuous dynamics of irregular time-distributed domains, which cannot be handled by existing TDG methods designed for discrete-dynamics and fixed-interval times.
@@
To address all the challenges, we propose the Koopman operator-driven continuous temporal domain generalization framework (Koodos). Specifically, the Koodos framework articulates the evolutionary continuity of the predictive models in CTDG, then leverages a continuous dynamic approach to model its smooth evolution over time. Koodos further simplifies the nonlinear model system by projecting them into a linearized space via the Koopman Theory. Finally, Koodos provides an interface that reveals the internal model dynamic characteristics, as well as incorporates prior knowledge and constraints directly into the joint learning process.
2 Related works
@@
Domain Generalization (DG) and Domain Adaptation (DA). DG approaches attempt to learn a model from multiple source domains that generalize well to an unseen domain [38; 37; 28; 5; 13; 56]. Existing DG methods can be classified into three strategies [49]: (1) Data manipulation techniques, such as data augmentation [45; 46; 57] and data generation [32; 40]; (2) Representation learning focuses on extracting domain-invariant features [17; 18] and disentangling domain-shared from domain-specific features [30]; (3) Learning strategies encompass ensemble learning [34], meta-learning [26; 14; 9], and gradient-based approaches [21]. Unlike DG, DA methods require simultaneously accessing source and target domain data to facilitate alignment and adaptation [50; 51; 29]. The technique includes domain-invariant learning [17; 47; 35; 48; 53], domain mapping [6; 20; 15; 31], ensemble methods [43], and so on. Both DG and DA are limited to considering generalization across categorical domains, which treats domains as individuals but ignores the smooth evolution of them over time.
@@
Temporal Domain Generalization (TDG). TDG is an emerging field that extends traditional DG techniques to address challenges associated with time-varying data distributions. TDG decouples time from the domain and constructs domain sequences to capture its evolutionary relationships. S-MLDG [27] pioneers a sequential domain DG framework based on meta-learning. Gradient Interpolation (GI) [39] proposes to extrapolate the generalized model by supervising the first-order Taylor expansion of the learned function. LSSAE [41] deploys a probabilistic framework to explore the underlying structure in the latent space of predictive models. DRAIN [2] constructs a recurrent neural network that dynamically generates model parameters to adapt to changing domains. TKNets [57] minimize the divergence between forecasted and actual domain data distributions to capture temporal patterns.
@@
Despite these studies, traditional TDG methods are limited to requiring domains presented in discrete time, which disrupts the inherent continuity of changes in data distribution, and the generalization can only be carried forward by very limited steps. No work treats time as a continuous variable, thereby failing to capture the full dynamics of evolving domains and generalize to any moment in the future.
@@
Continuous Dynamical Systems (CDS). CDS are fundamental in understanding how systems evolve without the constraints of discrete intervals. They are the study of the dynamics for systems defined by differential equations. The linear multistep method or the Runge-Kutta method can solve the Order Differential Equations (ODEs) [19]. Distinguishing from traditional methods, Neural ODEs [11] represent a significant advancement in the field of CDS. It defines a hidden state as a solution to the ODEs initial-value problem, and parameterizes the derivatives of the hidden state using a neural network. The hidden state can then be evaluated at any desired time using a numerical ODEs solver. Many recent studies have proposed on which to learn differential equations from data [42; 22; 23; 36].
@@ 3 Problem definition
Unlike traditional temporal domain generalization approaches [39; 2; 52; 57] that divide time into discrete intervals and require domain data be collected at fixed time steps, the CTDG problem treats time as continuous variable, allowing for the time points in training and test set to be any arbitrary positive real numbers, and requiring models to deal with the temporal domains as continuous streams. CTDG problem poses several unprecedented, substantial challenges in: 1) Characterizing the continuous dynamics of data and how that decides model dynamics; 2) Modeling the low-dimensional continuous dynamics of the predictive model, which are embedded in high-dimensional space due to over-parametrization; and 3) Optimizing and analyzing the continuous predictive model system, including the application of inductive biases to control its behavior.
@@ 4 Methodology
Figure 2: Macro-flows and micro-constraints in the proposed model framework.
@@ 4.1 Characterizing the continuous dynamics of the data and the model
Here, $\mathcal{L}$ symbolizes the loss function tailored to the prediction task. By employing Eq. 7, the model dynamics are defined for all observation time, and $\phi$ can be obtained by optimizing Eq. 5 and Eq. 6 via gradient descent. After that, the parameters of the predictive model can be calculated at any specific time using ODE solvers.
@@ 4.2 Modeling nonlinear model dynamics by Koopman operators
The aforementioned learning approaches are, however, limited in efficient modeling and prediction due to the entangled nonlinear dynamics in the high-dimensional parameters space. Identifying strongly nonlinear dynamics through such space is particularly susceptible to overfitting [8], leading to future predictive models eventually diverging from bifurcation to chaos. Despite the complexity of $h$, it is governed by the data dynamics $f$, which are typically low-dimensional and exhibit simple, predictable patterns. This suggests that the governing equations for $h$ are sparse within the over-parameterized space, and thus allow for a more simplified and manageable representation in a properly transformed space. Motivated by this, we propose leveraging the well-established Koopman Theory [24; 7] to represent these complex parameter dynamics. Koopman Theory provides a method for the global linearization of any nonlinear dynamics. It expresses the complex dynamic system as an infinite-dimensional Koopman operator acting on the Hilbert space of the system state measurement functions, in which the nonlinear dynamics will become linearized.
@@ 4.3 Joint optimization of model and its dynamics with prior knowledge
We introduce a comprehensive optimization approach designed to ensure that the system accurately captures the dynamics of the data. This process requires the joint optimization of several interconnected components under the optional constraint of additional prior knowledge about the underlying dynamics: the predictive model parameters $\theta_{1:T}$, the transformation functions $\varphi$ and $\varphi^{-1}$, and the Koopman operator $\mathcal{K}$. Our primary objectives are threefold: 1) Ensuring high prediction accuracy, 2) Maintaining consistency of parameters across different representations and transformations, and 3) Learning the Koopman invariant subspaces effectively. Fig. 2(b) illustrates the role of each constraint within our system: we manage two sets of states, intrinsic and integral, aligning across three spaces.
Predictive Model Parameters ($\theta$): Each observation time corresponds to a predictive model, which is tasked with making predictions and serving as the initial values for the dynamical system. They are primarily optimized to minimize the prediction error $L_{intri}$ of different domains:
$L_{intri}=\sum\nolimits_{i=1}^{T}\mathcal{L}(Y_{i},g(X_{i};\theta_{i})).$ (10)
Koopman Network Parameters ($\varphi$, $\varphi^{-1}$, $\mathcal{K}$): We estimate a function that transforms the parameters of the predictive model into Koopman space. Meanwhile, these predictive model parameters must be stable and consistent when converted between representations. This is realized by three constraints:
@@
    Reconstruction Loss: An autoencoder is leveraged to map parameter space to Koopman space by encoder $\varphi$ and back to the original space by decoder $\varphi^{-1}$. $L_{recon}$ ensures consistency between $\theta$ and its reconstructed form via transformations:
@@
    Dynamic Fidelity Loss: This term ensures that the transformation produces a Koopman invariant subspaces in which the dynamics can forward correctly. It measures the fit of the Koopman operator $\mathcal{K}$ against the transferred parameter:
@@
    Consistency Loss: It measures the consistency between the original and the dynamic parameter in the model space:
@@
Additionally, we load the dynamically integral parameters $\theta^{j\rightarrow i}_{i}$ back into the predictive model to evaluate its predictive capability, quantified by $L_{integ}=\sum_{i=1}^{T}\sum_{j=1}^{i}\mathcal{L}(Y_{i},g(X_{i};\theta^{j\rightarrow i}_{i}))$. Finally, the system optimizes the following combined loss to refine all components simultaneously:
$\{\theta_{1:T},\varphi,\varphi^{-1},\mathcal{K}\}=\mathop{\arg\!min}_{\theta_{1:T},\varphi,\varphi^{-1},\mathcal{K}}(\alpha L_{intri}+\alpha L_{integ}+\beta L_{recon}+\gamma L_{dyna}+\beta L_{consis}),$ (14)
with $\alpha$, $\beta$, and $\gamma$ as adjustable weights to balance the magnitude of each term, ensuring that no single term dominates during training.
During inference, given the moment $t_{s}$, the model uses the state from the nearest observation moment $t_{obs}$ as an initial value, integrating over the time interval $[t_{obs},t_{s}]$ in Koopman space to give the generalized model state $\theta^{obs\rightarrow s}_{s}$ at the desired test moment.
Analysis and Control of the Temporal Domain Generalization. Integrating Koopman theory into continuous temporal domain modeling facilitates the application of optimization, estimation, and control techniques, particularly through the spectral properties of the Koopman operator. We remark that the Koopman operator serves as a pivotal interface for analyzing and controlling the generalization process. Constraints imposed on the Koopman space will be equivalently mapped to the Model space. For instance, the eigenvalues of $\mathcal{K}$ are crucial as they provide insights into the system’s stability and dynamics, as illustrated below:
@@
    System Assessment. The generalization process is considered stable if all $\lambda_{i}$ satisfy $\text{Re}(\lambda_{i})<0$. Conversely, the presence of any $\lambda_{i}$ such that $\text{Re}(\lambda_{i})>0$ indicates instability in the system. When $\text{Re}(\lambda_{i})=0$, the system may exhibit oscillatory behavior. By analyzing the locations of these poles in the complex plane, we can assess the system’s long-term dynamics, helping us identify whether the generalized model is likely to collapse in the future.
@@
    Behavioral Constraints. Adding explicit constraints to $\mathcal{K}$ can guide the generalization toward desired behaviors. This process not only facilitates the incorporation of prior knowledge about domains but also tailors the system to specific characteristics. To name just a few, if the data dynamics are known to be periodic, configuring $\mathcal{K}$ as $\mathcal{K}=B-B^{T}$, with $B$ as learnable parameters, ensures that the model system exhibits consistent periodic behavior since the eigenvalues of $B-B^{T}$ are purely imaginary values. Additionally, employing a low-rank approximation such as $\mathcal{K}=UV^{T}$, with $U,V\in\mathbb{R}^{n\times k}$ and $k<n$, allows the model to concentrate on the most significant dynamical features and explicitly control the degrees of freedom of the generalization process.
@@
Theorem 2 (Superiority of continuous-time methods over discrete-time methods (informal)).
Continuous-time methods have smaller or equal errors compared to discrete-time methods in approximating temporal distribution drift, due to its consideration of the integral of changes over time.
The formal version and proof of Theorem 2 are given in Appendix C. We also provide a detailed model complexity analysis in Appendix A.1.4.